# MARBLE-Net: Learning to Localize in Multipath Environment with Adaptive Rainbow Beams

### v1

@@ IV-D Qualitative and Error Distribution Analysis
To gain deeper insight into the system’s behavior, we visualize the spatial error distribution of the MARBLE-Net (adaptive) model in Fig. 4. The visualization aligns perfectly with the quantitative results in Table II. The ‘L’ scene (Fig. 4d) shows the lowest error, with the vast majority of the area in dark blue (low error). The ‘rounded-L’ scene (Fig. 4c) also shows excellent performance, appearing visually superior to the ‘LoS-only’ scene (Fig. 4a), which has larger green and yellow high-error regions. In stark contrast, the ‘circle’ scene (Fig. 4b) exhibits large swathes of high-error (yellow, orange, and red) zones. These high-error zones correspond to areas where the uni-directional multipath effects are most severe, creating signal ambiguities that even the adaptive MARBLE-Net struggles to resolve completely. This visualization clearly illustrates that structured multipath (‘L’, ‘rounded-L’) is exploited, while uni-directional multipath (‘circle’) remains a significant challenge.
The cumulative distribution function (CDF) of localization error, shown in Fig. 5, offers additional insight into performance consistency across environments. The ‘L’ scene exhibits the most favorable distribution, with the curve rising steeply and reaching about 95% probability at a $2\text{\,}\mathrm{m}$ error threshold. The ‘rounded-L’ scene also performs well, clearly outperforming the ‘LoS-only’ case. In contrast, the ‘circle’ scene shows a long-tailed distribution due to unidirectional ambiguous multipath, with large errors not fully displayed in the figure. These observations are consistent with Table II, confirming that structured multipath is beneficial, whereas uni-directional multipath remains detrimental.
@@ IV-E Impact of Transmit Power
Figure 6: Localization RMSE vs. UAV Transmit Power for MARBLE-Net (adaptive).
We evaluate the system’s robustness to noise by varying the UAV transmit power, as shown in Fig. 6. As expected, the localization RMSE improves for all scenes with increased transmit power due to the higher SNR. Notably, the performance hierarchy across scenes remains consistent, with the ‘L’ and ‘rounded-L’ scenes outperforming the ‘LoS-only’ case, and the ‘circle’ scene performing the worst. This confirms that the geometric impact of multipath (whether beneficial or detrimental) is consistent across different noise levels.
@@ IV-F Complexity and Latency Analysis
RaiNet [klus2025deep] 1.95 M params $0.0180\text{\,}\mathrm{m}\mathrm{s}$
@@
Table III compares computational complexity and latency, measured on an NVIDIA RTX 5880 GPU with dual AMD EPYC 7Y43 CPUs. Although MARBLE-Net has far more parameters (13.96 M vs. 1.95 M for RaiNet), its average inference latency ($0.0172\text{\,}\mathrm{m}\mathrm{s}$) remains comparable to RaiNet ($0.0180\text{\,}\mathrm{m}\mathrm{s}$). Both models achieve over two orders of magnitude faster inference than the k-NN baseline ($2.21\text{\,}\mathrm{m}\mathrm{s}$), which depends on a 484 MB codebook and costly neighbor search. These results confirm the proposed model’s balance between accuracy and efficiency, enabling real-time deployment.
@@ V Conclusion
This paper has introduced MARBLE-Net, an end-to-end deep learning framework that jointly optimizes an adaptive rainbow beamformer and a localization network for ISAC systems. By leveraging ray-tracing-generated datasets across environments of varying complexity, we have demonstrated that a data-driven approach can improve localization performance. Our results show that co-designing the sensing beam with the neural network processor, guided by a structured multi-stage training strategy, allows the system to learn to exploit multipath information, reducing localization error from $5.85\text{\,}\mathrm{m}$ and $1.88\text{\,}\mathrm{m}$ to $0.88\text{\,}\mathrm{m}$ in a multipath-rich scene. This work validates that for high-precision localization in complex environments, the joint design of the physical-layer sensing strategy and the higher-layer processing network is not merely beneficial, but essential. Our findings also highlight a nuanced interaction with the propagation environment: while confined uni-directional multipath can hinder performance, we show that structured and directional multipath can be actively exploited to achieve accuracy surpassing even LoS-only conditions.
Future work will extend this framework to 3D and multi-target localization, and explore online adaptation in dynamic environments.

### v3

@@ IV-D Qualitative and Error Distribution Analysis
To gain deeper insight into the system’s behavior, Fig. 4 visualizes the spatial error distribution of the MARBLE-Net (adaptive) model, which is consistent with the quantitative results in Table II. The ‘L’ scene (Fig. 4d) exhibits the best performance, with most regions characterized by low localization errors. The ‘rounded-L’ scene (Fig. 4c) shows strong overall performance but reveals spatial heterogeneity: the upper region near the semi-circular arc suffers from larger errors due to curved geometry inducing uni-directional reflections with limited angular spread, whereas the lower region near the straight wall achieves higher precision, as planar surfaces provide structured and directional specular reflections that can be effectively exploited as virtual anchors. This comparison highlights the nuanced role of multipath propagation: structured, directional multipath from planar geometries serves as a valuable localization resource, potentially outperforming LoS-only conditions, while confined uni-directional multipath from curved structures introduces ambiguities that degrade performance.
The cumulative distribution function (CDF) of localization error in Fig. 5 provides further insight into performance consistency across different environments. The ‘L’ scene exhibits the most favorable distribution, with a steep rise reaching nearly 100% probability at a $2\text{\,}\mathrm{m}$ error threshold, indicating high localization stability. The LoS-only case follows as the second-best scenario, while the ‘rounded-L’ and ‘circle’ scenes show more gradual curves, with the ‘rounded-L’ performance lying between the two. The ‘circle’ scene exhibits the least favorable distribution, reflecting the difficulty of resolving signal ambiguities in uni-directional multipath environments. These trends are consistent with the results in Table II, confirming that structured multipath is highly beneficial for localization, whereas the geometry of the ‘circle’ scene remains a major limitation.
@@ IV-E Impact of Transmit Power
We evaluate the system’s robustness to noise by varying the UAV transmit power, as shown in Fig. 6. As expected, the localization RMSE improves for all scenes with increased transmit power due to the higher SNR. Notably, the performance hierarchy across scenes remains consistent, with the ‘L’ scene consistently outperforming the LoS-only case, while the ‘rounded-L’ and ‘circle’ scenes exhibit higher errors due to their respective multipath complexities. This confirms that the geometric impact of multipath (whether beneficial or detrimental) is consistent across different noise levels.
@@ IV-F Complexity and Latency Analysis
RaiNet [10] 1.95 M params $0.0180\text{\,}\mathrm{m}\mathrm{s}$
@@
Table III compares computational complexity and latency, measured on an NVIDIA RTX 5880 GPU with dual AMD EPYC 7Y43 CPUs. Although MARBLE-Net has significantly more parameters than RaiNet (13.96 M vs. 1.95 M), its average inference latency ($0.0172\text{\,}\mathrm{m}\mathrm{s}$) is marginally lower ($0.0180\text{\,}\mathrm{m}\mathrm{s}$). This result is attributed to the GPU’s ability to efficiently parallelize the wider layers in MARBLE-Net, combined with the use of the efficient LeakyReLU activation, as opposed to the more expensive Tanh function used in RaiNet. Both models achieve over two orders of magnitude faster inference than the k-NN baseline ($2.21\text{\,}\mathrm{m}\mathrm{s}$), which depends on a 484 MB codebook and costly neighbor search. These results confirm the proposed model’s balance between accuracy and efficiency, enabling real-time deployment.
@@ V Conclusion
This paper has introduced MARBLE-Net, an end-to-end deep learning framework that jointly optimizes an adaptive rainbow beamformer and a localization network for ISAC systems. By leveraging ray-tracing-generated datasets across environments of varying complexity, we have demonstrated that a data-driven approach can improve localization performance. Our results show that co-designing the sensing beam with the neural network processor, guided by a structured multi-stage training strategy, allows the system to learn to exploit multipath information, reducing localization error from $5.85\text{\,}\mathrm{m}$ and $1.33\text{\,}\mathrm{m}$ to $0.61\text{\,}\mathrm{m}$ in a multipath-rich scene. This work validates that for high-precision localization in complex environments, the joint design of the physical-layer sensing strategy and the higher-layer processing network is not merely beneficial, but essential. Our findings also highlight a nuanced interaction with the propagation environment: while confined uni-directional multipath can hinder performance, we show that highly structured and directional multipath, as observed in the ‘L’ scene, can be actively exploited to achieve accuracy surpassing even LoS-only conditions, turning multipath from an impairment into a localization resource.
Future work will extend the framework to 3D multi-target localization and adaptation in dynamic environments.